OpenStudy (anonymous):

which is greater 3 4/5 or 3 2/3

TheSmartOne (thesmartone):

They both have 3 infront of it. So do you think 4/5 is greater or less than 2/3

TheSmartOne (thesmartone):

Remember 4/5=.8 and 2/3=.666 Which one is greater? @hannahmckinney

OpenStudy (anonymous):

put it this way... 4/5 is .8 or 80% of a whole, but 2/3 is equal to .66666 or 66% of a whole. you have 3 wholes. adding the percentages we get 380%= 3 4/5 or 366%= 3 2/3. So 3 4/5 is larger than 3 2/3. Do you understand my approach?

TheSmartOne (thesmartone):

You don't have to put it in to percentages...

OpenStudy (anonymous):

Yes i know, but if she understands it that way better; then it's a better way to teach her. :)